# A Guide to Misinformation Detection Datasets

### v1

@@ A.5 Implementation Details of GPT-4 with Web Search Predictive System
A.6 Contradiction Between Ground Truth Label and Predictive System
@@
A.7 Supplement on Manual Labeling of Prediction Validity
@@
A.8 Supplement on Contradiction Evaluator
@@
Now that youve ready the statement and assessments, rate how much the assessments contradict or not on a scale from 0 (no contradiction) to 10 (complete contradiction). However, you must not state your score until youve presented a concise analysis. Do not begin your response with a number. First write your analysis, then write a vertical bar "|", then finally state your contradiction score.
@@
Now that youve ready the statement and assessments, answer if the assessments contradict or not. However, you must not state your decision until youve presented a concise analysis. Do not begin your response with a label. First write your analysis, then write a vertical bar "|", then finally "1: contradiction" or "0: no contradiction".
@@
Now that youve ready the statement and assessments, answer if the assessments contradict or not. However, you must not state your decision until youve presented a concise analysis. Do not begin your response with a label. First write your analysis, then write a vertical bar "|", then finally "1: contradiction" or "0: no contradiction", or if you are not sure write "-1: unsure".
@@
Appendix B Note on Computational and Other Resources
All political leaning, baselines, and contradiction evaluator experiments were run using the OpenAI API, with the web search version of the latter drawing on Cohere web connector in an agentic setup (Appendix A.5). The total cost was approximately \$1,500 USD. Other analyses were run on authors’ computers or Google Colab, requiring minimal resources and expenditure.

### v2

@@ C.1. Implementation Details of GPT-4 with Web Search Predictive System
C.2. Contradiction Between Ground Truth Label and Predictive System
@@
C.3. Supplement on Manual Labeling of Prediction Validity
@@
C.4. Supplement on Contradiction Evaluator
@@
Now that youve ready the statement and assessments, rate how much the assessments contradict or not on a scale from 0 (no contradiction) to 10 (complete contradiction). However, you must not state your score until youve presented a concise analysis. Do not begin your response with a number. First write your analysis, then write a vertical bar ”|”, then finally state your contradiction score.
@@
Now that youve ready the statement and assessments, answer if the assessments contradict or not. However, you must not state your decision until youve presented a concise analysis. Do not begin your response with a label. First write your analysis, then write a vertical bar ”|”, then finally ”1: contradiction or ”0: no contradiction”.
@@
Now that youve ready the statement and assessments, answer if the assessments contradict or not. However, you must not state your decision until youve presented a concise analysis. Do not begin your response with a label. First write your analysis, then write a vertical bar ”|”, then finally ”1: contradiction or ”0: no contradiction”, or if you are not sure write ”-1: unsure”.